Two cars leave towns 800 kilometers apart at the same time and travel toward each other. One car's rate is 12 kilometers per hour more that the other's. If they meet in 5 hours, what is the rate of the faster car?
Do not do any rounding.

Answer by richwmiller(17219)

800=5*x+5*(x+12)
x=74
x+12=86 kmph for the faster car
check
800=5*74+5*86
800=370+430
800=800
ok